OpenStudy (campayne):

Use elimination to solve the following system of equations. {-x + y + 2z = 7 2x + 3y + z = 1 -3x - 4y + z = 4 x = 1, y = 2, z = -2 x = 2, y = 1, z = -2 x = -2, y = 1, z = 2 x = 1, y = -2, z = -2

OpenStudy (mathmale):

Hope the following suggestions will be of help: -x + y + 2z = 7 Mult. this equation by 2 and add your result to the second equation. 2x + 3y + z = 1 show your result. -3x - 4y + z = 4

OpenStudy (mathmale):

-x + y + 2z = 7 Next, mult this equation by -3 and combine your result with the 2x + 3y + z = 1 third equation. Show your result. -3x - 4y + z = 4 By doing these things you should be able to eliminate the variable x.

OpenStudy (campayne):

-2 1 and 2 right?